# Multi-Model Framework for Reconstructing Gamma-Ray Burst Light Curves

### v1

@@ 3 Results
We compute the reduction in the percentage of the error fractions in order to evaluate the enhancement in fit post reconstruction.
%$\%_{DEC}=\frac{\left|EF^{\rm{after}}_{X}\right|-\left|EF^{\rm{before}}_{X}\right|}{\left|EF^{\rm{before}}_{X}\right|}\times 100.$ (25)
The reconstruction for each category of GRBs for all models is shown in Fig. 8, and the histogram distribution of the relative percentage decrease for the three parameters is illustrated in Fig. 9.
Table 3 shows how effectively the models’ reconstructed data performs in comparison to the initial observed data for all GRBs. Although the models perform well for most GRB reconstructions, a few outliers are observed in the W07 parameters. We classify outliers as instances where the relative percentage decrease exceeds 100%.
@@
Figure 9: Distribution plot of all three W07 parameters in a grid with parameters (left to right): i) $\log F_{a}$ (Column 1) ii) $\log T_{a}$ (Column 2) iii) $\alpha$ (Column 3) and the models (top to bottom): i) Deep GP (Row 1); ii) TCN Model (Row 2); iii) CNN-BiLSTM Model (Row 3); iv) BNN Model (Row 4); v) Polynomial Curve Fitting Model (Row 4); vi) Isotonic Regression Model (Row 6)
GRB ID $EF_{\log_{10}(T_{i})}$ $EF_{\log_{10}(F_{i})}$ $EF_{\alpha_{i}}$ $EF_{\log_{10}(T_{i})}$ RC $EF_{\log_{10}(F_{i})}$ RC $EF_{\alpha_{i}}$ RC %${}_{\log_{10}(T_{i})}$${}_{\log_{10}(F_{i})}$${}_{\alpha_{i}}$
@@
% log10(Ta) % log10(Fa) % $\alpha$ logTa % logFa % $\alpha$
@@
TCN -22.6 -28.1 -38.7 13.2 11.6 2.20
DGP -31.7 -32.6 -41.0 7.24 6.86 1.48
CNN-LSTM -32.2 -32.8 -39.5 2.57 2.94 0.550
BNN -20.2 -25.6 -33.4 4.04 4.04 0.550
Polynomial Curve Fitting -32.3 -33.1 -41.9 2.57 2.57 0.920
Isotonic Regression -36.2 -36.1 -43.6 3.12 3.30 0.730
Bi-Mamba -33.3 -33.6 -41.9 2.70 2.70 0.700
MLP -25.9 -28.6 -37.7 3.60 3.30 0.900
Bi-LSTM -28.4 -28.2 -36.0 3.80 3.60 1.20
GP (W07) -22.8 -23.2 -33.5 3.10 3.30 1.10
W07 model (10%) -24.5 -25.7 -36.2 2.40 2.40 0.900
W07 model (20%) -21.2 -22.9 -34.7 2.80 2.80 1.50
218 Good GRBs Dainotti et al. (2023c)
TCN -20.2 -26.2 -37.8 14.6 10.5 1.37
DGP -35.4 -36.2 -45.4 3.25 3.25 0.460
CNN-LSTM -31.7 -32.7 -41.0 0.450 0.450 0
BNN -17.2 -24.2 -32.4 2.75 2.29 0.450
Polynomial Curve Fitting -30.3 -32.5 -42.1 0.450 0.450 0.450
Isotonic Regression -38.8 -37.4 -46.8 0.450 0.450 0
Bi-Mamba -32.7 -34.3 -43.6 0.400 0.400 0.400
MLP -24.1 -27.8 -38.9 0.900 0.400 0.400
Bi-LSTM -26.1 -28.0 -37.3 1.30 1.80 1.30
GP (W07) -25.6 -27.9 -41.6 0.400 0 0
W07 model (10%) -33.3 -35.0 -43.3 0 0 0
W07 model (20%) -29.5 -31.2 -40.6 0 0 0
Table 3: Summary of average uncertainty decrease and outliers for different reconstruction models on 545 GRBs and the 218 good GRBs subset.
@@ 4 Summary and Conclusion
    The isotonic regression model achieved strong reductions in uncertainty - $36.3\%$ for $\log T_{a}$, $36.1\%$ for $\log F_{a}$, and $43.6\%$ for $\alpha$. These reductions are among the highest, particularly for $\alpha$, where it outperformed several deep learning models. Isotonic regression remains a satisfactory baseline, particularly strong for the $alpha$ parameter. This model achieved a 3.80% improvement for the $\alpha$ parameter over the Bi-Mamba model.
@@
    The TCN model achieved moderate improvements in $\log T_{a}$ and $\log F_{a}$ parameters, but showed a notable reduction of 38.7% in the uncertainty of the $\alpha$ parameter.
@@
    For the $\alpha$ parameter specifically, TCN performs slightly better than the MLP Manchanda et al. 2024, with a 2.60% improvement. However, it exhibits the highest number of outliers among all models. This suggests that although TCN performs well on seen data, it may struggle with generalization, particularly when applied to gaps in GRB LC.
@@
    The DGP model effectively models the uncertainties in the data. It performs notably well for the 218 good GRBs, achieving 31.7% reduction in $\log(T_{a})$, 32.6% in $\log(F_{a})$, and 41.0% in $\alpha$. It also maintains outlier percentages of 7.24% ($\log(T_{a})$), 6.86% ($\log(F_{a})$), and 1.48% ($\alpha$), indicating reliable generalization performance. Its performance is also comparable to the Bi-Mamba model, showcasing similar uncertainty values for $\log(T_{a})$, $\log(F_{a})$, and $\alpha$ parameters.
@@
    On comparison with the GP model used in Dainotti et al. 2023c; Manchanda et al. 2024 for the 545 sample, we observe that the DGP model performs better across all three uncertainty reduction parameters, highlighting the importance of introducing layers that are able to capture complex patterns that a single-layer GP fails to catch. However, there is a notable increase in the outliers. The differences are demonstrated in Table 4.
@@
    Comparing the DGP model with the hybrid model GP-RF mentioned in Dainotti et al. 2023c; Manchanda et al. 2024 for the entire sample, shows that the DGP model outperforms the hybrid model as well. Although the outliers are slightly higher than GP-RF, the DGP model reduces the outliers for the $\alpha$ parameters slightly. The comparison is shown in Table 4.
  • 7.
    The CNN-LSTM model demonstrates moderate yet consistent performance for all three parameters. It achieves reductions of 32.2%, 32.8%, and 39.5% for $\log(T_{a})$, $\log(F_{a})$, and $\alpha$ respectively. Notably, it also shows low outlier rates (2.57% and 2.94% for $\log(T_{a})$ and $\log(F_{a})$), and the lowest outlier percentage for $\alpha$ at just 0.550%, indicating reliability.
  • 8.
    Compared with the LSTM model used in Manchanda et al. 2024, the CNN-LSTM model outperforms it across all three parameters and shows fewer outliers. This highlights the benefit of combining convolutional layers with recurrent structures to effectively capture both local and temporal features in GRB light curves, see Table 5 for a detailed comparison.
  • 9.
    The BNN model reduces the W07 parameters by the lowest percentage for the full dataset as well as the 218 Good GRBs. It achieves reductions of 20.2%, 25.6%, and 33.4% for $\log(T_{a})$, $\log(F_{a})$, and $\alpha$ respectively. It did show the lowest outlier for the $\alpha$ parameter of 0.550
  • 10.
    Polynomial Curve Fitting demonstrates solid performance in uncertainty reduction, achieving 32.3%, 33.1%, and 41.9% decreases in $\log(T_{a})$, $\log(F_{a})$, and $\alpha$, respectively. It retains low outlier rates, particularly 2.57% in both $\log(T_{a})$ and $\log(F_{a})$, and 0.920% in $\alpha$. It is on par with the Bi-Mamba model, with an equal decrease of 41.9% in the $\alpha$ parameter.
@@
    We consider these $logT_{a}$, $logF_{a}$, and $\alpha$ as the most important outputs of our reconstruction. The parameter $\alpha$ describes the post-plateau decay behavior, relevant for the testing of the standard fireball model (Piran, 1999), and the reduction of $\alpha$ parameter is crucial to obtain reduced uncertainties on the closure relationship (Dainotti et al., 2021, 2024a). The parameters $\log T_{a}$ and $\log F_{a}$ are key for identifying empirical correlations, such as the Dainotti relations in two and three dimensions (Dainotti et al., 2008, 2010, 2011, 2013, 2016, 2020a, 2020b, 2022b; Levine et al., 2022). Among all models tested, Isotonic Regression demonstrates the most significant reduction in error for the three parameters. Thus, we recommend it as the preferred model when accurate estimation of these physically meaningful quantities is the main objective.
    Parameters DGP GP % Improvement GP-RF % Improvement
    % Uncertainty Decrease
    $\log F_{a}$31.7 -22.8 39.0 -27.5 15.3
    $\log T_{a}$32.6 -23.2 40.5 -25.0 30.4
    $\alpha$41.0 -33.5 22.4 -34.7 18.2
    % Outliers
    $\log F_{a}$ 7.24 3.10 -134 5.30 -36.6
    $\log T_{a}$ 6.86 3.30 -108 5.60 -22.5
    $\alpha$ 1.48 1.10 -25.7 2.30 55.4
    Table 4: Comparison between DGP & GP and DGP & GP-RF
    Parameters CNN-LSTM Bi-LSTM % Improvement
    % Uncertainity Decrease
    $\log T_{a}$32.2 -28.4 13.4
    $\log F_{a}$32.8 -28.2 16.3
    $\alpha$39.5 -36.0 9.72
    % Outliers
    $\log T_{a}$ 2.57 3.80 32.4
    $\log F_{a}$ 2.94 3.60 18.3
    $\alpha$ 0.550 1.20 54.2
    Table 5: Comparison between CNN-LSTM and Bi-LSTM
@@
These reconstructions achieve a significant reduction in uncertainty, which improves the reliability of the GRB plateau parameters. By achieving reduced uncertainty, the relationships associated with plateau emissions can be considered as standard candles for cosmological research (Dainotti et al., 2022g, c, 2023b).
@@
These methods can also be explored in future studies, including cutting-edge deep learning approaches such as neural ordinary differential equations (neural ODEs), Physics Informed Neural Network (PINN), and evolutionary algorithms. Additionally, the potential of transformer architectures and large language models to capture complex temporal patterns warrants investigation. When combined with new observational data, it is anticipated that these methods will greatly improve LCR’s accuracy, dependability, and interpretability.
@@
Though we have only used Swift LCs to develop this reconstruction framework thus far, we plan to apply it to other ongoing missions like SVOM (Atteia et al., 2022) and Einstein Probe (Yuan et al., 2022), as well as upcoming missions like THESEUS (Amati et al., 2018) and HiZ-GUNDAM (Yonetoku et al., 2024). Additionally, future analysis will provide data from other wavelengths. Given the availability of the most comprehensive optical catalogue to date, it is particularly appealing to extend this work to optical wavelengths (Dainotti et al., 2020a, 2022b, 2024a, 2024b; Zhang and Mészáros, 2004).
5 Appendix
@@
This section discusses an additional model we experimented with, namely the Time-aware Neural Ordinary Differential Equation (ODE). Although this model did not yield satisfactory outcomes for our particular application, we believe it may prove more beneficial in other hybrid configurations or in future research endeavors.
5.1 Time-aware Neural ODE (TN-ODE)
@@
We developed the TN-ODE model to handle time-series data in which the measurements are irregular, as is frequently the case with GRB light curves. The model consists of three main components:
@@
    Encoder: After reading the supplied data, this module compresses it into a representation with fewer dimensions. It uses a time-aware LSTM network (Baytas et al., 2017), a variant of RNNs that incorporates the time gaps between observations to control how much past information is retained or forgotten.
@@
    Neural ODE Block: The compressed latent representation from the encoder is then passed into a Neural ODE block (Chen et al., 2018). This component models the continuous-time dynamics of the latent state, simulating how it evolves over time. The neural ODE enables interpolation and extrapolation at arbitrary time points, even between known data samples (Rubanova et al., 2019).
@@
To incorporate uncertainty and improve generalization, we employ variational sampling (Kingma et al., 2013) during encoding. Stochasticity is introduced into the latent vectors via sampling from a training-learned distribution. This step allows the model to generate a distribution over possible outputs, enabling it to produce not only point forecasts but also confidence intervals, which quantify the model’s uncertainty.
@@
In our experiments, TN-ODE demonstrated strong performance in modeling and forecasting GRB LCs with irregular time intervals. However, due to the model’s complexity and large number of parameters, careful hyperparameter tuning was required to prevent overfitting and mitigate sensitivity to noise. While the uncertainty estimates were generally informative, they were sometimes unstable because of the stochastic sampling process.

### v2

@@ 2.6 Parameter Uncertainty and Evaluation Metrics
We compute the reduction in the percentage of the error fractions in order to evaluate the enhancement in fit post-reconstruction.
$\%_{DEC}=\frac{\left|EF^{\rm{after}}_{X}\right|-\left|EF^{\rm{before}}_{X}\right|}{\left|EF^{\rm{before}}_{X}\right|}\times 100.$ (30)
3 Results
The reconstruction for each category of GRBs for all models is shown in Fig. 9, and the histogram distribution of the relative percentage decrease for the three parameters is illustrated in Fig. 10 and tabulated in Table 2.
Table 3 shows how effectively the models’ reconstructed data perform in comparison to the initial observed data for all GRBs. Although the models perform well for most GRB reconstructions, a few outliers are observed in the W07 parameters. We classify outliers as instances where the relative percentage increase exceeds 100%. Table 4 shows the performance strengths and weaknesses of each model.
@@ 3 Results
Figure 10: Distribution plot of all three W07 parameters in a grid with parameters (left to right): i) $\log F_{a}$ (Column 1) ii) $\log T_{a}$ (Column 2) iii) $\alpha$ (Column 3) and the models (top to bottom): i) Deep GP (Row 1); ii) TCN Model (Row 2); iii) CNN-BiLSTM Model (Row 3); iv) BNN Model (Row 4); v) Polynomial Curve Fitting Model (Row 5); vi) Isotonic Regression Model (Row 6); vii) QSS Model (Row 7)
GRB ID $EF_{\log_{10}(T_{i})}$ $EF_{\log_{10}(F_{i})}$ $EF_{\alpha_{i}}$ $EF_{\log_{10}(T_{i})}$ RC $EF_{\log_{10}(F_{i})}$ RC $EF_{\alpha_{i}}$ RC $_{\log_{10}(T_{i})}$ $_{\log_{10}(F_{i})}$ $_{\alpha_{i}}$
@@
% log10(Ta) % log10(Fa) % $\alpha$ % logTa % logFa % $\alpha$
@@
QSS -43.5 -43.2 -48.3 2.69 3.84 0.960
Polynomial Curve Fitting -20.8 -21.6 -27.4 2.88 2.30 1.15
CNN-BiLSTM -20.3 -20.9 -25.1 3.07 2.69 0.768
Isotonic Regression -18.0 -18.5 -24.0 3.07 2.88 0.960
BNN -10.9 -17.6 -15.9 5.95 5.76 2.69
DGP -11.6 -12.3 -15.9 6.91 5.95 1.92
TCN -5.31 -12.7 -16.2 18.5 14.6 4.62
Attention U-Net -37.9 -38.5 -41.4 1.73 2.50 1.34
MLP -37.2 -38.0 -41.2 1.73 2.30 1.34
GP (W07) -16.9 -18.6 -24.3 3.07 3.45 1.54
W07 model (10%) -18.0 -19.1 -25.2 2.30 2.30 2.11
W07 model (20%) -15.8 -17.8 -23.8 2.30 2.69 2.30
207 Good GRBs (dainotti2023stochastic)
QSS -48.0 -48.8 -55.1 0 0.483 0
CNN-BiLSTM -23.1 -25.2 -29.1 1.93 1.93 0
Polynomial Curve Fitting -21.0 -24.0 -30.2 0 0 0
Isotonic Regression -19.4 -21.6 -28.0 0.966 0.966 0.483
DGP -14.8 -16.5 -20.7 4.35 3.86 0.966
BNN -11.2 -19.7 -19.1 2.42 2.90 0.483
TCN -3.56 -16.4 -16.9 21.7 18.8 4.83
Attention U-Net -38.8 -40.3 -44.0 0.483 0.483 0
MLP -38.1 -39.3 -43.9 0.483 0.483 0
GP (W07) -17.2 -20.0 -28.8 1.93 1.45 0
W07 model (10%) -23.0 -24.9 -30.8 0.966 0.966 0
W07 model (20%) -21.2 -23.2 -28.9 0 0 0
Table 3: Summary of average uncertainty decrease and outliers for different reconstruction models on 521 GRBs and the 207 good GRBs subset.
Attributes QSS Polynomial Curve Fitting CNN-BiLSTM DGP BNN TCN Isotonic Regression
High Uncertainty Reduction
Good Sparse Data Performance
Almost Zero Outliers for Good GRBs
Low Computational Cost
Captures Complex Patterns
Table 4: Comparison of the models presented here on various attributes.
@@ 4 Summary and Conclusion
    The QSS model achieves outstanding reductions in uncertainty: 43.5% for $\log T_{a}$, 43.2% for $\log F_{a}$, and 48.3% for $\alpha$. QSS’s reduction is almost double the reductions obtained by the other five models presented in this article. These reductions are the lowest we have achieved, particularly for $\alpha$, where it outperformed several deep learning models. For Good GRBs, it has strong reductions of 48.0%, 48.8%, and 55.1% for $\log(T_{a})$, $\log(F_{a})$, and $\alpha$ respectively. It maintains zero outliers across $\log(T_{a})$ and $\alpha$ parameters for Good GRBs. QSS remains a satisfactory baseline, particularly strong for the $\alpha$ parameter.
@@
    The Polynomial Curve Fitting model achieves good reductions in uncertainty - $20.8\%$ for $\log T_{a}$, $21.6\%$ for $\log F_{a}$, and $27.4\%$ for $\alpha$. These reductions are the highest among our newly introduced models, particularly for $\alpha$, where it outperformed several deep learning models. It maintains zero outliers across all three parameters for Good GRBs. Polynomial Curve Fitting model is the fourth top-performing baseline, among all the models we have tried so far. This model’s performance is comparable to the Bi-Mamba model.
@@
    The CNN-BiLSTM model achieves moderate performance across all parameters, 20.3%, 20.9%, and 25.1% for $\log(T_{a})$, $\log(F_{a})$, and $\alpha$ respectively. It maintains zero outliers for the $\alpha$ parameter for Good GRBs. This model shows good reduction across all GRB parameters with the lowest outlier rate for $\alpha$ (0.768%). Similarly, the DGP model demonstrates uncertainty reduction of $11.6\%$ for $\log T_{a}$, $12.6\%$ for $\log F_{a}$, and $15.9\%$ for $\alpha$ parameters.
@@
    The Isotonic Regression models lowers uncertainty by $18.0\%$ for $\log T_{a}$, $18.5\%$ for $\log F_{a}$, and $24.0\%$ for $\alpha$. The BNN model showcases moderate reductions of 10.9%, 17.6%, and 15.9% for $\log(T_{a})$, $\log(F_{a})$, and $\alpha$ respectively. While the TCN model achieves reductions of $5.31\%$ for $\log T_{a}$, $12.7\%$ for $\log F_{a}$, and $16.2\%$ for $\alpha$.
@@
    Most of the outliers found in the study are GRBs which have at least two of the variations–flares, bumps, breaks, and double breaks. Very few Good GRBs (0-4% across the three parameters) are present in the outliers, as demonstrated in the table featuring Good GRB outlier rates in the paper. This concludes that the models perform better on Good GRBs as compared to the other three types.
@@
    The platinum sample, composed of 50 GRBs from dainotti2020a is benefitting the most from this analysis and it will be then used for theoretical model discrimination and as a cosmological tool with the advantage of having a decreased uncertainties on the cosmological parameters as it was discussed in dainotti2022g where a reduction of 47.5% on the uncertainties parameters allow to reach the precision of SNe Ia in (Conley2011) already now and the one reached by (Betoule) in 17 years from now. So this reduced percentages allow to reach these goals in a few years less.
@@
    We consider these $logT_{a}$, $logF_{a}$, and $\alpha$ as the most important outputs of our reconstruction. The parameter $\alpha$ describes the post-plateau decay behavior, relevant for the testing of the standard fireball model (piran1999gamma), and the reduction of the $\alpha$ parameter is crucial to obtain reduced uncertainties on the closure relationship (dainotti2021closure; dainotti2024analysis). The parameters $\log T_{a}$ and $\log F_{a}$ are key for identifying empirical correlations, such as the Dainotti relations in two and three dimensions (Dainotti2008; dainotti2010a; dainotti2011a; dainotti2013determination; dainotti2016fundamental; dainotti2020a; dainotti2020b; dainotti2022b; levine2022). Among all the models tested, QSS demonstrates the most significant reduction in error for all three parameters. Comparing it with our previous models, QSS persisted as the most effective reconstruction strategy. Thus, we recommend QSS as the preferred model when accurate estimation of these physically meaningful quantities is the main objective.
@@
These reconstructions achieve a significant reduction in uncertainty, which improves the reliability of the GRB plateau parameters for both model estimator and for their use as standard candles. By achieving reduced uncertainty, the relationships associated with plateau emissions can be considered as standard candles for cosmological research (dainotti2022g; dainotti2022c; dainotti2023b). Additionally, the precision in determining cosmological parameters has also been improved. The results of this work can approach the estimation accuracy of dark matter from Type Ia supernovae (dainotti2022g; Betoule).
@@
These methods can also be explored in future studies, including cutting-edge deep learning approaches such as neural ordinary differential equations (neural ODEs), Physics Informed Neural Network (PINN), and evolutionary algorithms. We analysed a time-series model; for details, refer to Appendix 5. Additionally, the potential of transformer architectures and large language models to capture complex temporal patterns warrants investigation. When combined with new observational data, it is anticipated that these methods will greatly improve LCR’s accuracy and interpretability.
While significant uncertainty reduction was achieved on the parameters, the challenge of providing global influence in relation to GRBs as distributions still remains. Our current models exhibit limited global influence during parameter tuning and do not account when GRBs are considered altogether, because we analyze GRBs singularly. The models used in this analysis primarily function as curve-fitters with deep local influence, inferring/extracting local representations to fill LC gaps. The shared hyperparameters across GRBs contribute to a shared influence between GRBs; however, this influence is limited and warrants further exploration. Optimizing the hyperparameters for 16 GRBs, taken from a set of 4 GRBs belonging to each class (Good, Flares/Bumps, Breaks, Flares with bumps + double breaks) and freezing these hyperparameters for all the 521 GRBs provides a first step towards a global framework for reconstructing the samples that is not GRB specific. Future work could build on this methodology to address sparsity and obtain consistent time-length samples. These reconstructed samples could then be fed into temporal aware ML/ DL methods to extract parameters across several GRBs. This will allow us to include global GRB parameters in the latent features, rather than fine tuning singular GRBs.
@@
Though we have only used Swift LCs to develop this reconstruction framework thus far, we plan to apply it to other ongoing missions like SVOM (atteia2022svom) and Einstein Probe (yuan2022einstein), as well as upcoming missions like THESEUS (amati2018theseus) and HiZ-GUNDAM (yonetoku2024high). Additionally, future analysis will provide data from other wavelengths. Given the availability of the most comprehensive optical catalogue to date, it is particularly appealing to extend this work to optical wavelengths (dainotti2020optical; dainotti2022b; dainotti2024analysis; dainotti2024optical).
5 Appendix A. TN-ODE
@@
This section discusses an additional model we analyzed with, namely the Time-aware Neural Ordinary Differential Equation (ODE). Although this model did not yield satisfactory outcomes for our particular application, we believe it may prove more beneficial in other hybrid configurations or in future research endeavors.
@@
We developed the TN-ODE model to handle time-series data in which the measurements are irregular, as is frequently the case with GRB LCs. The model consists of three main components:
@@
    Encoder: After reading the supplied data, this module compresses it into a representation with fewer dimensions. It uses a time-aware LSTM network (baytas2017patient), a variant of RNNs that incorporates the time gaps between observations to control how much past information is retained or forgotten.
@@
    Neural ODE Block: The compressed latent representation from the encoder is then passed into a Neural ODE block (chen2018neural). This component models the continuous-time dynamics of the latent state, simulating how it evolves over time. The neural ODE enables interpolation and extrapolation at arbitrary time points, even between known data samples (rubanova2019latent).
@@
To incorporate uncertainty and improve generalization, we employ variational sampling (kingma2013auto) during encoding. Stochasticity is introduced into the latent vectors via sampling from a training-learned distribution. This step allows the model to generate a distribution over possible outputs, enabling it to produce not only point forecasts but also confidence intervals, which quantify the model’s uncertainty.
@@
In our analysis, TN-ODE demonstrated strong performance in modeling and forecasting GRB LCs with irregular time intervals. However, due to the model’s complexity and large number of parameters, careful hyperparameter tuning was required to prevent overfitting and mitigate sensitivity to noise. While the uncertainty estimates were generally informative, they were sometimes unstable because of the stochastic sampling process.
6 Appendix B. Shift Problem Resolution
In this section, we detail an investigation into systematic shifts at larger times in the reconstructed $\log T_{a}$ and $\log F_{a}$ parameters, relative to the original GRBs (see Figure 11). The systematic shift in the platinum sample is shown in Figure 13. This issue was persistent across all models utilized in this study. This deviation consequently produced an elevated scatter value in the reconstructed platinum sample, and a different slope of the 2-D Dainotti relation between the rest-frame end time $T_{a}$ and its luminosity, $L_{a}$. a result that is contradictory to the research’s main goal.
To resolve this anomaly, the following investigative methods were employed:
  • 1.
    Noise Recalibration: The reconstruction’s noise generation protocol was recalibrated. Instead of assuming uniform or Gaussian distributions, samples were drawn from the best-fit flux error distribution. This change was implemented to mitigate artificial bias in the post-$T_{a}$ regions. However, it was observed that the noise level was not the cause of the systematic shift.
  • 2.
    Hybrid Model Testing: We explored multiple hybrid models, such as interfacing an ML model with a power-law decay curve. A separate strategy applied a double-power law to the ML model output following a GRB-specific threshold. None of these hybrid attempts, however, produced a significant reduction in the observed shift or the overall scatter.
  • 3.
    $\sigma$-Cut Filtering: A filtering mechanism using a $\pm 2\sigma$ cut was applied to the $(\log T_{a},\log F_{a})$ distribution. The purpose was to eliminate extreme outliers that were presumed to be artifacts of the reconstruction process. Also, this cut did not resolve the issue.
  • 4.
    Simulation Testing: We simulated synthetic GRBs with predefined $\log T_{a}$, $\log F_{a}$, and $\alpha$ parameters. By introducing controlled data gaps of varying magnitudes, we analyzed the conditions that produced the shift. The analysis revealed that the shift was predominantly caused by large flux gaps located immediately after the $T_{a}$ point.
  • 5.
    Data-Augmentation and gap-aware reconstruction: To address this core issue, we developed a ‘gap-aware’ reconstruction algorithm. This method first identifies significant data gaps (defined as $\Delta\log t>0.05$). It then adaptively inserts new data points exclusively into these identified sparse regions. The quantity of new points is determined adaptively as a percentage of the original LC’s length: 5% (for >500>500 points), 10% (for 250–500 points), 30% (for 100–250 points), and 40% (for <100<100 points, minimum 20). This ensures that larger gaps are filled with proportionally more data, enhancing continuity without oversampling the denser segments.
The gap-aware reconstruction method succeeded in resolving the systematic shift (Figure 12) and was then applied across the complete 521 GRB sample. After implementing these corrections, the systematic shift was gone (Figure 14), and the reconstructed GRBs achieved a scatter ($\sigma_{\text{BiMamba}}=0.5134$) statistically comparable to the original platinum sample ($\sigma_{\text{Platinum}}=0.5131$), remaining consistent within $1\sigma$.
Figure 11: Systematic shift observed in the platinum sample post Bi-MAMBA Model reconstruction. Units: $\log F$ (erg cm -2s-1) and $\log T$ (s).
Figure 12: Resolved shift post Bi-MAMBA Model reconstruction. Units: $\log F$ (erg cm -2s-1) and $\log T$ (s).
Figure 13: Systematic shift observed in the platinum sample post BiMamba Model reconstruction for $\log L_{a}$ (erg s-1) and $\log T_{a}$ (s).
Figure 14: Resolved shift post gap-aware BiMamba Model reconstruction for $\log L_{a}$ (erg s-1) and $\log T_{a}$ (s).
CRediT authorship contribution statement
A. Kaushal: Conceptualization, Methodology, Supervision, Project Administration, Investigation, Software, Writing Original draft - Review & Editing. A. Manchanda: Conceptualization, Methodology, Supervision, Project Administration, Investigation, Software, Writing Original draft - Review & Editing. M. G. Dainotti: Conceptualization, Methodology, Supervision, Project Administration, Software, Resources, Writing Original draft - Review & Editing. Krishnanjan Sil: Conceptualization, Methodology, Software, Validation, Formal Analysis, Investigation, Data Curation, Writing - Original Draft, Writing - Review & Editing, Visualization. K. Gupta: Conceptualization, Methodology, Software, Validation, Formal Analysis, Investigation, Writing - Original Draft, Writing - Review & Editing. Z. Nogala: Methodology, Software, Validation. A. Madhan: Methodology, Software, Validation. S. Naqi: Methodology, Software, Validation. R. Kumar: Methodology, Software, Validation, Formal analysis, Optimization, writing -OriginalDraft. V. Oad: Methodology, Software, Validation. N. Indoriya: Methodology, Software, Validation. D.H. Hartmann: Writing - Review & Editing. M. Bogdan: Methodology. A. Pollo: Methodology. N. Fraija: Writing - Review & Editing. J.X. Prochaska: Methodology. N. Fraija: Methodology.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
Data will be made available upon reasonable request to the corresponding author.